Weekly Inspection- June 12th

Observations:

Weather has continued to be very warm – temperatures now averaging mid 20Cs. No rain in sight for the next two weeks. Forage starting to get less abundant but still flowering:

green alkanet, elderflower, wisteria, pyracantha and geranium just starting.

All colonies have continued to expand and as yet no signs of swarming. Starting to see much more stores being collected.

Hive Record
Colony
Q
BIAS
Brood
Stores
CommentNext Inspection
AquaNY 5 Frames of bias with expanded pattern Quite of lot nectar, will add super Add QE and super Laying up drone trap - uncap in 2 weeks
BlueYY9 frames of bias - still plenty of laying space 2/3rd of 1st super and starting to work 2ndStill expanding with lots of eggsCheck on drone trap
DenimYY6 frames of bias- so has expanded Started working the super- small amount of nectar Removed drone trap- 3.5 per 100 - good result check on super activity